OpenStudy (anonymous):

y=1/5x-6 complete the ordered pairs (5, )(0, )

OpenStudy (immanuelv):

plug in 5 for x and 0 for x for your answer :D

OpenStudy (anonymous):

yeah the fraction confuses me

OpenStudy (immanuelv):

is the question \[y = \frac{ 1 }{ 5 } x -6?\]

OpenStudy (anonymous):

yes.

OpenStudy (immanuelv):

so 5 * 1/5 = 1 :) and 0 * 1/5 = 0 :) does that help?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

do i multiply the numerator only?

OpenStudy (immanuelv):

\[\frac{ 1 }{ 5 } (5) - 6 \] \[1-6 =?\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

-5

OpenStudy (immanuelv):

yup -5 ... thats your answer for (5, -5)

OpenStudy (immanuelv):

the second one is @ x= 0 so when you multiply (1/5) with 0 you get 0. then you do 0-6 =?

OpenStudy (anonymous):

so basically the 5 cancels out?

OpenStudy (immanuelv):

for the first one since the numerator and denominator is 5

OpenStudy (immanuelv):

for fractions, you multiply the numerator and divide the denominator for ex (1/5)*5 = 1*5 = 5 then you do 5/5 = 1 :D

OpenStudy (anonymous):

ooooh ok thank you!!!!!! :)

OpenStudy (anonymous):

y=1/5x-6 when x=5 y=1/5*5-6=1-6=-5 (5,-5) when x=0 y=1/5*0-6=0-6=6 (0,-6)

OpenStudy (immanuelv):

you're welcome :D

OpenStudy (nincompoop):

what a poor way of explaining things

OpenStudy (nincompoop):

you're just basically giving out answers without teaching proper concepts